Diffusion Morning Report 10/10/06

Ted Berg tberg at stanford.edu
Tue Oct 10 06:14:10 PDT 2006


Good Morning All,
    Looking OK this morning. All LPs still down until tank fix is 
completed. LTO tube change in progress. Epi still has slightly high 
base- still looking into it.
Burn box minora changed yesterday. No other problems at this time.

Diffusion Mike,Ray, Ted 	  	  	  	  	 
  	Date: 	10/10/06 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	Comments 	  	  	 
Tylan 1 	Up 	  	  	  	 
Tylan 2 	Up 	  	  	  	 
Tylan 3 	Up 	  	  	  	 
Tylan 4 	Up 	  	  	  	 
Tylan 5-BBr3 	Up 	  	  	  	 
Tylan 6-POCl3 	Up 	  	  	  	 
Tylan 7-TEOS2 	Up 	  	8.41/30 	  	 
Tylan 8-SiGe 	Up 	  	6.72/30 	  	 
Tylan 9-Poly 	Up 	  	.61/30 	  	 
Tylan 10-Nitride 	Up 	  	8.52/30 	  	 
Tylan 11-BPSG-LTO 	Down 	pull and clean 	0/30 	  	 
Tylan 12-FGA 	Up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	Up 	  	  	  	 
Thermco 2 	Up 	  	  	  	 
Thermco Poly 	Up 	  	  	  	 
Thermco Nitride 	Up 	  	  	  	 
Epi 	Down 	high base pressure 	  	  	 
  	
	
	
	
	
Burn Box 	up 	  	  	  	 
ZYGO 	Up 	  	  	  	 
  	up 	  	   	  	 
GAAS 22 	  	  	  	  	 
GAAS 23 	UP 	  	  	  	 
SiH4 for furnaces 	up 	  	  	  	 

	  	B on line 260 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20061010/85e76636/attachment.html>


More information about the diffusion mailing list